} likr st = 65uf ON ___ory 94+ -°F ATN Pvt
Answer #1

Please find the solution in the images below

Image 1

Initially the Charged fully switch is closed the capacitor and acts as open circuit o Voltage across the capacitor at t w 33K

for Explanation please refer the below images

Image 2

Discharge Phase : . . upto now we discussed about the charging phase of a Capacitior. At any point the swild hoint the switch

Image 3

In this case the capacitor acts as battery and reverse current direction and discharges completty. g dir the current R ww et

Image 4

(t-o) = loge Velt) - loge E e che = loge Velt) loga-log 6 = log a -t/RC . EL e = Volt) -t/Rc. [VeLE)= Ee the .. (((t)= c d Ve

Thank you

Have a nice day